OpenStudy (anonymous):

Help! I need steps! a line passes through (-7,-5) and (-5,4) a. write an equation for the line in point-slope form b. rewrite the equation in standard form using integers

OpenStudy (anonymous):

Please help!!!

OpenStudy (anonymous):

(-7,-5) and (-5,4) can you find the slope from these points???

OpenStudy (whpalmer4):

Equation for the slope of a line passing through points \((x_1,y_1), (x_2,y_2)\): \[m = \frac{y_2-y_1}{x_2-x_1}\] Point-slope form for a line passing through point \((x_0,y_0)\) with slope \(m\) is \[y-y_0 = m(x-x_0)\] Standard form is \(Ax+Bx = C\), \( A\ne0,B\ne0\)